# Classifier-free graph diffusion for molecular property targeting

### v1

@@ 3 Classifier-Free Graph Diffusion Process
Figure 1: (a) The overall architecture of the proposed Neural Network. Notice that the guide $\bm{y}$ is now an input to the model. (b) The graph transformer layer for the input $\bm{X}$ (it is applied analogously to E and $\bm{u}$). (c) The self attention layer within the graph transformer layer. Notice that we superimposed ${}^{\prime}$ to the vectors indicate that they are network/layer outputs.
3.1 Classifier-free DiGress (FreeGress)
The forward noise process for our CF method can be straightforwardly inherited by discrete DDPM in general, and for the sake of this work from DiGress, as detailed in (1) and (2). The key impact of CF guidance in instead on the adaptive reverse process. To this end, we begin by rewriting the conditioned reverse process $p_{\theta}(G^{t-1}|G^{t},\bm{y})$ to incorporate $\bm{y}$ as follows:
$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y}),$ (8)
which requires to reformulate the marginals from Eq. 6. Specifically, we can rewrite the nodes marginal along the lines of Gu et al. (2022) as:
$\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y}),$ (9)
where $f_{\theta}$ is a neural network that predicts the true node types $\bm{x}^{0}$ using both the current noisy graph $G^{t}$ and the guide $\bm{y}$. Similarly, the edges marginal is expressed as:
$\sum_{\bm{e}\in\mathcal{E}}q(\bm{e}_{ij}^{t-1}|\bm{e}_{ij}^{t},\bm{e}_{ij}^{0}=\bm{e})\,f_{\theta}(\bm{e}_{ij}^{0}=\bm{e}|G^{t},\bm{y}).$ (10)
At this point, we have, in principle, a classifier-free DDPM. However, conditioning might not work as expected in this scenario, since the model could learn to completely ignore the guide $\bm{y}$. To prevent this undesirable behaviour, we leverage conditional dropout (Ho and Salimans 2022), where the conditioning vector $\bm{y}$ is replaced, with probability $p_{uncond}$, by a non-parametrized placeholder $\bm{\bar{y}}$. To summarize, the final vector $\bm{y}$ employed in the equations (and, consequently, in the neural network) is chosen as follows:
$\bm{y}\leftarrow\begin{cases}\bm{\bar{y}}&\text{with probability }p_{uncond}\\ \bm{y}&\text{with probability }1-p_{uncond}\end{cases}$ (11)
Intuitively, the placeholder signals the network to perform inference without guidance vector, as if the process was unconditioned. As a result, the model becomes capable of performing both conditioned and unconditioned sampling.
The major difference between (Ho and Salimans 2022) and our work is the fact that the former is designed for continuous signals, such as images, and their neural network is trained to predict the Gaussian noise $\epsilon$ corrupting the input signal $\bm{x}^{t}$ given both the latter and the conditioning vector $\bm{y}$. FreeGress, instead, is designed for discrete graphs and its neural network is trained to infer $G^{0}$ given the corrupted graph $\bm{G}^{t}$ and conditioning vector $\bm{y}$. To complete derivation of FreeGress, we build on Tang et al. (2023) and express $f_{\theta}$ as:
$f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y})=p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})\,p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})^{s},$ (12)
where the first term is the standard reverse process, and the second forces the model to take the guide into account, with $s\in\mathbb{R}_{+}$ being its relative importance. This same objective can be expressed in log-probability space as:
$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})}+(s+1)\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t}).$ (13)
Lastly, using Bayes’ rule, the term $\log p_{\theta}(\bm{y}|\bm{x}^{0}_{i},G^{t})$ is decomposed as:
$\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p_{\theta}(\bm{x}^{0}_{i}|G^{t})},$ (14)
which is computable since the model can output the unconditioned probability using the placeholder in place of the guide or, more formally, since ${p(\bm{x}^{0}_{i}|G^{t})}\approx{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{\bar{y}}})$. The process is symmetric for the edges $\bm{e}\in\mathcal{E}$. Finally, to add more flexibility, we parametrize $\bm{\bar{y}}$ so that it is learned along with the parameters $\theta$.
Neural Architecture
@@ 3.1 Classifier-free DiGress (FreeGress)
In order to accommodate the explicit guidance, we modified the original DiGress architecture as shown in Figure 0(a). Basically, the network consists of a stack of Graph Transformer layers. In its original implementation, DiGress allows to augment the input matrices $\bm{X}$ and E with arbitrary node and edge features deemed useful (indicated with the $+$ superscript in Figure 0(a)). In addition, a vector $\bm{u}$ containing the current time step $t$ is added, which itself be enriched with graph-wise features. All these data are augmented with the guide $\bm{y}$ and processed by the stack of graph transformer layers. The graph transformer is displayed in Figure 0(b): at a high level, it resembles a standard transformer architecture, with a self attention module followed by dropout, residual connection, layer norm, and a feedforward stack. The self attention layer (shown in Figure 0(c)) first processes the (augmented) $\bm{X}$ matrix through the standard outer product and scaling typical of attention layers. Then, the result is used in conjunction with the edge tensor E as input to a FiLM layer (Perez et al. 2018) to produce unnormalized attention scores. The scores are passed through a softmax layer, combined once again with the values $\bm{X}$. The attention vectors are then flattened and mixed with the vector $\bm{u}$ through another FiLM layer, and then transformed linearly to obtain the prediction logits. The vector $\bm{u}$ is instead processed differently, as shown visually in the rightmost part of Figure 0(c). Specifically, $\bm{X}$ and E are first passed through independent PNA layers (Corso et al. 2020) and then summed together with a linear projection of $\bm{u}$ to form the final representation.
Loss function.
@@
$\delta\sum_{1\leq i\leq n}\mathrm{NLL}\left(\bm{x}_{i}^{0},f_{\theta}(\bm{x}_{i}^{t})\right)+\gamma\sum_{1\leq i,j\leq n}\mathrm{NLL}\left(\bm{e}_{ij}^{0},f_{\theta}(\bm{e}_{ij}^{t})\right),$
which is the negative log-likelihood (NLL) between the actual node (resp. edge) types and those predicted by the neural network described above. Parameters $\delta$ and $\gamma$ adjust the preference towards nodes or edges prediction.
Conditioned Inference
At each reverse process iteration, FreeGress generates a slightly denoised graph by first computing from the network $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{y})}$, the conditioned log probability, and $\log{p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{\bar{y}})}$, the unconditioned one. The final probability is obtained by exponentiating the sum of the two terms while giving an arbitrary (usually larger) weight $s$ to the conditioned probability:
$\log{p(\bm{x}^{0}_{i}|G^{t})}+(s+1)\,(\log{p(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p(\bm{x}^{0}_{i}|G^{t})}),$
where we omit exponentiation for conciseness, which corresponds to $f_{\theta}(\bm{x}^{0}=\bm{x}|G^{t},\bm{y})$ in Eq. 12. Then, the posteriors $q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})$ are computed $\forall\bm{x}\in\mathcal{X}$ and summed as in Eq. 9. This procedure is executed for each atom in the graph, and the resulting probabilities are sampled to obtain the nodes of the graph $G^{t-1}$. The process is symmetric for the graph edges.
3.2 Graph size inference
One limitation of diffusion models as a whole is the fact that the size (for graphs specifically, the number of nodes) of a sample cannot change during the denoising process. Some studies consider the possibility of inserting and removing elements from a generated, mono-dimensional sequence (Johnson et al. 2021) but, to the best of our knowledge, there are no similar works for graphs. In the original implementation of DiGress, the number of nodes the generated graph will have is sampled from the marginal distribution computed from the training and validation sets, respectively. While this is not an issue for unconditioned sampling, we observed that for conditioned sampling, a certain property (e.g. molecular weight) might be featured only by molecules with a specific number of atoms. To solve this problem, we propose that the number of nodes $n$ of the graph to be generated is sampled from $p_{\varepsilon}(n|\bm{y})$, which is parametrized as a neural network with two hidden layers and a softmax output layer. Here, the idea is to exploit the guide even before generation starts, by providing a graph size which is correlated with the requested property.
@@ 4 Experiments
Here, we detail the experimental analysis to evaluate FreeGress on property targeting tasks, that is, generating compounds which meet pre-specified molecular properties.
4.1 Dataset
Our experiments were conducted on QM9, a dataset of 133k small molecules (up to 9 heavy atoms) which was also part of the evaluation of DiGress, and ZINC-250k, a collection of 250k drug-like molecules selected from the ZINC dataset. For the latter, the molecules were first preprocessed by removing non-neutral charges and stereochemistry information, reducing the dataset size to 204k molecules.
@@ 4.2 Metrics and targets
All the experiments were performed using the following scheme. After we trained each model, we randomly sampled 100 molecules from the dataset and computed the desired target properties as vectors $\bm{y}$, which were used as conditioning vectors. Then, we used the vectors 10 times each to perform conditioned generation, for a total of 1000 generated molecules for each model. We computed the desired properties on the generated samples [^1] Using packages such as RDKit (Landrum 2016) and psi4 (Turney et al. 2012) and compared the results with the properties of the original samples. The metric chosen for the comparison is the Mean Absolute Error (MAE) between the target properties and the properties of the generated molecules:
@@
where $\bm{y}_{i}$ are the target properties of the $i$-th molecule from the dataset and $\widehat{\bm{y}}_{i,j}$ is the target property of the $j$-th molecule generated using the properties of the $i$-th molecule as guide. The targeted properties were dipole moment $\mu$ and the Highest Occupied Molecular Orbital (HOMO) for QM9; for ZINC-250k, we targeted log-partition coefficient (LogP) and Molecular Weight (MW).
Model Hyper-parameters Extra features HOMO $\downarrow$ $\bm{\mu}$downarrow HOMO + $\bm{\mu}$downarrow
DiGress Not provided No $0.56\pm 0.01$ $0.81\pm 0.04$ $0.87\pm 0.03$
FreeGress $p_{uncond}=0,s=5$ No $0.50\pm 0.05$ $0.99\pm 0.08$ $0.74\pm 0.06$
$p_{uncond}=0,s=5$ Yes $0.34\pm 0.03$ $0.95\pm 0.07$ $0.62\pm 0.04$
$p_{uncond}=0.1,s=5$ No $0.30\pm 0.04$ $0.60\pm 0.06$ $0.59\pm 0.07$
$p_{uncond}=0.1,s=5$ Yes $\mathbf{0.28\pm 0.04}$ $\mathbf{0.54\pm 0.06}$ $\mathbf{0.53\pm 0.07}$
$p_{uncond}=0.2,s=5$ No $0.31\pm 0.05$ $0.64\pm 0.07$ $0.62\pm 0.06$
$p_{uncond}=0.2,s=5$ Yes $0.36\pm 0.10$ $0.57\pm 0.07$ $0.55\pm 0.06$
Model Hyper-parameters LogP MW
MAE $\downarrow$ Valid $\uparrow$ MAE $\downarrow$ Valid $\uparrow$
DiGress $\lambda=50$ $1.15\pm 0.09$ 75% $73.40\pm 7.10$ 75%
$\lambda=500$ $0.60\pm 0.05$ 68% $60.30\pm 6.10$ 73%
$\lambda=1000$ $0.49\pm 0.05$ 65% $61.10\pm 5.90$ 69%
$\lambda=3000$ $0.52\pm 0.13$ 39% $64.10\pm 5.90$ 59%
FreeGress $p_{uncond}=0.1,s=2$ $\mathbf{0.15\pm 0.02}$ 85% $15.73\pm 2.74$ 58%
$p_{uncond}=0.1,s=2$, NI $0.16\pm 0.01$ 84% $\mathbf{15.18\pm 2.71}$ 75%
Table 1: Results on the QM9 (top) and ZINC-250k (bottom) datasets. Property targeting has been evaluated using Mean Absolute Error (see Section 4.2). HOMO: Highest Occupied Molecular Orbital; $\mathbf{\mu}$: dipole moment; $\lambda$: weight of the classier gradient for CB guidance; $s$: weight of conditioned inference for FreeGress; $p_{uncond}$: conditional dropout parameter for FreeGress; Valid: percentage of valid molecules among those sampled; NI: node inference (see Section 3.2).
@@ 4.3 Experimental details
On QM9, we trained a set of FreeGress variants against the published results of DiGress. In the original article, DiGress was trained for the property targeting task without additional features, and hyper-parameters such as $\lambda$ (which regulates the impact of the regressor’s gradient) were not provided. FreeGress instances were trained with $s=5$, and with $p_{uncond}=\{0,0.1,0.2\}$ to study the effect of conditional dropout on the generative process. Moreover, we also trained different variants with and without additional features for $\bm{X}$ and $\bm{u}$, to understand if they provided any positive impact. Additional graph-wise features considered for this study were number of $k$-cycles with $k\in\{1,\ldots,6\}$, number of connected components, first 5 non-zero eigenvalues, molecular weight; among node-wise features, we considered the atom’s valency, the number of $k$-cycles a node belongs to, and an estimation of the biggest connected component size. FreeGress used similar hyper-parameters (5 graph transformer layers; embedding size of 256 for $\bm{X}$ and 128 for both E and $\bm{u}$; 8 attention heads; $T=500$, 1200 training epochs with batch size 512; Amsgrad optimizer (Reddi, Kale, and Kumar 2018) with learning rate of $2^{-4}$ and weight decay of $1^{-12}$; $\gamma=2\delta$) as DiGress for a fair comparison. Importantly, since FreeGress does not require an auxiliary regressor, all the variants are trained with half the number of parameters than DiGress (4.6 million vs. 9.2 million, approximately). All models required approximately seven hours to train on a nVidia A100 with 80GB of VRAM.
ZINC-250k.
@@
For this dataset, we used a slightly different experimental setup since DiGress was not evaluated on this dataset previously. Specifically, we trained from scratch both DiGress and FreeGress with 12 layers and 1000 training epochs with batch size 256, while keeping the other hyper-parameters the same as for QM9. We used larger models since ZINC-250k includes bigger (in terms of number of atoms) and more diverse molecules. Training required approximately 100 hours for 1000 epochs each on the same machine as above. While FreeGress variants used approximately 16 million parameters, DiGress used approximately 32 million including the property regressor. We trained without additional features, since preliminary experiments have revealed that they had little to no influence on the performance, while at the same time largely increasing the training time. For DiGress, we evaluated different variants alternatives with $\lambda=\{50,500,1000,3000\}$. For FreeGress, we fixed $s=2$ and $p_{uncond}=0.1$ and evaluated one variant with the inference method proposed in Section 3.2, and one without. In the former case, $p_{\varepsilon}(n|\bm{y})$ was parameterized as a neural network with 2 layers with 512 units each, and ReLU activations.
Input $\mu$: 2.9326Estimated $\mu$: 2.8836
Input $\mu$: 4.1600Estimated $\mu$: 4.0071
Input HOMO: -6.7892Estimated HOMO: -6.6912
Input $\mu$: 2.9326Estimated $\mu$: 2.8836
Input $\mu$: 4.1600Estimated $\mu$: 4.0071
Input HOMO: -6.7892Estimated HOMO: -6.6912
Input HOMO: -7.4478Estimated HOMO: -7.3225
Figure 2: Curated molecules from the QM9 dataset.

### v2

@@ 3.0.2 Reverse Process
Figure 2: (a) The overall architecture of FreeGress. Notice that the guide $\bm{y}$ is an input to the model. (b) The graph transformer layer. $\bm{V}$ represents the triple $(\bm{X},\bm{\textsf{E}},\bm{u})$. (c) The self attention layer within the graph transformer layer. Notice that we superimposed to the vectors indicate that they are network/layer outputs.
3.0.3 Neural Architecture.
@@
Figure 2(a) depicts the neural architecture of our approach integrating explicit guidance. The network consists of a stack of Graph Transformer layers. The input matrices $\bm{X}$ and E can be augmented with arbitrary node and edge features deemed useful (indicated with the $+$ superscript in Figure 2(a)). In addition to the two matrices, a further vector $\bm{u}$ is used as input, containing the current timestep $t$ along with optional graph-level features (such as the overall molecular weight). All these data are augmented with the guidance $\bm{y}$ and processed by the stack of graph transformer layers. The graph transformer is displayed in Figure 2(b): at a high level, it resembles a standard transformer architecture, with a self attention module followed by dropout, residual connection, layer norm, and a feedforward stack. The self attention layer (shown in Figure 2(c)) first processes the (augmented) $\bm{X}$ matrix through the standard outer product and scaling typical of attention layers. Then, the result is used in conjunction with the edge tensor E as input to a FiLM layer [15] to produce unnormalized attention scores. The scores are passed through a softmax layer, combined once again with the values $\bm{X}$. The attention vectors are then flattened and mixed with the vector $\bm{u}$ through another FiLM layer, and then transformed linearly to obtain the prediction logits. The vector $\bm{u}$ is instead processed differently, as shown visually in the rightmost part of Figure 2(c). Specifically, $\bm{X}$ and E are first passed through independent PNA layers [3] and then summed together with a linear projection of $\bm{u}$ to form the final representation.
3.0.4 Loss function.
@@
$\sum_{1\leq i\leq n}\mathrm{CE}\left(\bm{x}_{i}^{0},f_{\theta}(\bm{x}_{i}^{t})\right)+\gamma\sum_{1\leq i,j\leq n}\mathrm{CE}\left(\bm{e}_{ij}^{0},f_{\theta}(\bm{e}_{ij}^{t})\right),$ (12)
which is the cross-entropy (CE) between the actual node (resp. edge) types and those predicted by the neural network described above. $\gamma$ is a hyper-parameter which adjusts the preference towards edges prediction. When Equation 11 is employed, CE is replaced by the negative log-likelihood loss.
3.0.5 Conditioned Inference.
At each reverse process iteration, FreeGress generates a slightly denoised graph by first computing from the network $p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{y})$, the conditioned probability, and $p_{\theta}(\bm{x}_{i}^{0}|\bm{x}_{i}^{t},\bm{\bar{y}})$, the unconditioned one. The final probability is obtained by summing the two terms while giving an arbitrary (usually larger) weight $s$ to the conditioned probability, using Equation 10. Then, the posteriors $q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})$ are computed $\forall\bm{x}\in\mathcal{X}$ and summed as in Equation 7. This procedure is executed for each atom in the graph, and the resulting probabilities are sampled to obtain the nodes of the graph $G^{t-1}$. The process is symmetric for the graph edges.
3.1 Conditioning on the number of nodes
A general limitation of diffusion models is the fact that the size of a sample (for graphs specifically, the number of nodes) cannot change during the denoising process. Some studies consider the possibility of inserting and removing elements from a generated, mono-dimensional sequence [12] but, to the best of our knowledge, there are no similar works for graphs. In the original implementation of DiGress, the number of nodes the generated graph will have is sampled from the marginal distribution computed from the training and validation sets, respectively. While this is not an issue for unconditioned sampling, we observed that for conditioned sampling, a certain property (e.g. molecular weight) might be featured only by molecules with a specific number of atoms. To solve this problem, we propose that the number of nodes $n$ of the graph to be generated is sampled from $p_{\varepsilon}(n|\bm{y})$, which is parametrized as a neural network with two hidden layers and a softmax output layer. Here, the idea is to exploit the guide even before generation starts, by providing a graph size which is correlated with the requested property.
@@ 4 Experiments
Here, we detail the experimental analysis to evaluate FreeGress on property targeting tasks by generating compounds that meet pre-specified properties.
4.1 Datasets
Our experiments were conducted on QM9, a dataset of 133k small molecules (up to 9 heavy atoms) which was also part of the evaluation of DiGress, and ZINC-250k, a collection of 250k drug-like molecules selected from the ZINC dataset. For the latter, the molecules were first preprocessed by removing stereochemistry information. We have also removed scarcely present non-neutrally-charged atoms, leaving only N+ and O-. The latter have been treated as standalone atom types instead of their neutrally-charged counterparts. The final dataset size was consequently reduced to 228k molecules. Further statistics about the datasets are provided in the Supplementary Section 1.
@@ 4.2 Metrics and targets
All the experiments were performed using the following setup. After we trained each model, we randomly sampled 100 molecules from the dataset and computed the desired target properties as vectors $\bm{y}$, which were used as conditioning vectors. Then, we used the vectors 10 times each to perform conditioned generation, for a total of 1000 generated molecules for each model. We computed the desired properties on the generated samples [^1] Using packages such as RDKit [13] and psi4 [23] and compared the results with the properties of the original samples. The metric chosen for the comparison is the Mean Absolute Error (MAE) between the target properties and the properties of the generated molecules:
@@
where $\bm{y}_{i}$ are the target properties of the $i$-th molecule from the dataset and $\widehat{\bm{y}}_{i,j}$ is the target property of the $j$-th molecule generated using the properties of the $i$-th molecule as guide. The targeted properties were the dipole moment $\mu$ and the Highest Occupied Molecular Orbital (HOMO) for QM9; for ZINC-250k, we targeted the Log-Partition coefficient (LogP), and the Quantitative Estimation of Drug-likeness (QED). The proposed node inference method was evaluated on the ZINC-250k dataset targeting Molecular Weight (MW).
@@ 4.3 Experimental details
We trained and compared a set of FreeGress variants against a set of DiGress models, including an unconditional variant as baseline. To the best of our knowledge, DiGress is the only model that allows for property targeting at the time of writing. FreeGress instances were trained with $s=\{1,3,5\}$, and with $\rho=\{0,0.1,0.2\}$ to study the effect of conditional dropout on the generative process. We have also experimented variants with additional features and $\rho=\{0,0.1\}$; in this case, $\rho=0.2$ was excluded as it did not improve on the former in preliminary trials. For DiGress, we evaluated the performance using $\lambda=\{200,400,600,800\}$. To ensure a fair comparison, we trained both DiGress and FreeGress with matching architectural design: specifically, we used 5 graph transformer layers; embedding size of 256 for $\bm{X}$ and 128 for both E and $\bm{u}$; 8 attention heads; $T=500$, 1200 training epochs with batch size 512; Amsgrad optimizer [18] with learning rate of $2^{-4}$ and weight decay of $1^{-12}$; $\gamma=2$. Importantly, since FreeGress does not require an auxiliary regressor, all the variants are trained with half the number of parameters than DiGress (4.6 million vs. 9.2 million, approximately). A single training required up to 12 hours on an nVidia V100 with 16GB of VRAM.
4.3.2 ZINC-250k for logP/QED targeting.
@@
For these experiments we used a slightly different setup. Specifically, we trained from scratch both DiGress and FreeGress with 12 layers, 1000 training epochs, batch size 256, and doubling the embedding size of $\bm{y}$, while keeping the other hyper-parameters the same. We used larger models since ZINC-250k includes bigger (in terms of number of atoms) and more diverse molecules. In this case, we did not perform experiments with additional features since they increased training time by a prohibitive amount. FreeGress variants required approximately 16 million parameters, while DiGress used approximately 32 million including the property regressor. Training the various models required approximately 5 days each on an nVidia A100 with 80GB of VRAM.
4.3.3 ZINC-250k for MW targeting.
These experiments evaluate the node inference method proposed in Section 3.1. As such, we targeted MW since it trivially depends on the number of nodes. The setup is similar to the logP/QED experiments, with some slight differences. In this case only, we have managed to train FreeGress with additional features and $\rho\in\{0,0.1\}$ ($\rho=0.2$ was excluded since it did not bring significant improvements in early trials). The node inference model $p_{\varepsilon}(n|\bm{y})$ was parameterized as a neural network with 2 layers with 512 units each and ReLU activations.
Table 1: Results on the QM9 dataset. First two columns are single conditioning, last column is multiple conditioning on $\mu$ and HOMO.
$\mu$ HOMO $\mu$+HOMO
MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
Unconditional $1.68\pm 0.15$ $91.5\%$ $0.95\pm 0.10$ $91.5\%$ $1.32\pm 0.1$ $91.5\%$
DiGress
$\lambda=200$ $1.06\pm 0.13$ $87.0\%$ $0.70\pm 0.07$ $91.7\%$ $1.13\pm 0.08$ $88.7\%$
$\lambda=400$ $0.95\pm 0.07$ $86.5\%$ $0.65\pm 0.07$ $91.7\%$ $1.08\pm 0.09$ $87.2\%$
$\lambda=600$ $0.80\pm 0.07$ $82.5\%$ $0.61\pm 0.07$ $91.2\%$ $1.11\pm 0.10$ $86.7\%$
$\lambda=800$ $0.86\pm 0.08$ $76.7\%$ $0.52\pm 0.05$ $90.5\%$ $1.12\pm 0.11$ $87.1\%$
FreeGress
$\rho=0$ $1.03\pm 0.10$ $86.7\%$ $0.45\pm 0.05$ $92.0\%$ $0.83\pm 0.11$ $82.3\%$
$\rho=0.1,s=1$ $1.06\pm 0.12$ $88.3\%$ $0.51\pm 0.07$ $92.3\%$ $0.87\pm 0.09$ $86.2\%$
$\rho=0.1,s=3$ $0.74\pm 0.08$ $83.7\%$ $0.32\pm 0.04$ $90.1\%$ $0.68\pm 0.09$ $83.4\%$
$\rho=0.1,s=5$ $0.67\pm 0.10$ $77.7\%$ $0.40\pm 0.08$ $92.9\%$ $0.77\pm 0.13$ $77.0\%$
$\rho=0.2,s=1$ $1.03\pm 0.10$ $89.1\%$ $0.47\pm 0.05$ $91.6\%$ $0.84\pm 0.10$ $84.3\%$
$\rho=0.2,s=3$ $0.83\pm 0.10$ $82.1\%$ $0.32\pm 0.04$ $93.8\%$ $0.80\pm 0.18$ $72.0\%$
$\rho=0.2,s=5$ $0.74\pm 0.11$ $75.5\%$ $0.44\pm 0.07$ $93.3\%$ $0.87\pm 0.19$ $72.7\%$
FreeGress+extra features
$\rho=0$ $0.90\pm 0.10$ $90.1\%$ $0.36\pm 0.04$ $94.2\%$ $0.65\pm 0.07$ $93.1\%$
$\rho=0.1,s=1$ $0.90\pm 0.10$ $90.0\%$ $0.35\pm 0.03$ $92.6\%$ $0.65\pm 0.07$ $92.5\%$
$\rho=0.1,s=3$ $0.60\pm 0.07$ $85.2\%$ $0.27\pm 0.04$ $92.1\%$ $0.54\pm 0.07$ $90.0\%$
$\rho=0.1,s=5$ $0.52\pm 0.07$ $81.5\%$ $0.40\pm 0.08$ $91.1\%$ $0.59\pm 0.11$ $84.9\%$